How many grams of Fe are produced by reacting 2.00 kg of Al with 300 g Fe2O3?


1
Expert's answer
2021-10-05T23:56:24-0400

2Al + Fe2O3 "\to" 2Fe +Al2O3


mass of Fe2O3 = 300/159.7

=1.8785


Mole ratio

Fe2O3 : Fe = 1 : 2


Moles of Fe = 2× 1.8785

= 3.757

Mass of Fe = 3.757 × 55.845


= 209.81g